OpenStudy (anonymous):

derivative of f(x)=8/x at x=-1

OpenStudy (anonymous):

Do you know how to solve derivatives?

OpenStudy (anonymous):

yes. I got the answer 8/x^2. But I'm not sure if it should be negative or not.

OpenStudy (amistre64):

how do you find your derivatives? limit rules or derivative shortcuts?

OpenStudy (anonymous):

I used difference quotient.

OpenStudy (amistre64):

then your application of it was slightly off

OpenStudy (anonymous):

I know the answer is going to be 8 or -8 but I'm not sure which one.

OpenStudy (amistre64):

\[\frac{x(8')-8(x')}{x^2}\] \[\frac{x(0)-8(1)}{x^2}\] \[\frac{-8}{x^2}\] now apply x=-1

OpenStudy (anonymous):

Okay, so the answer is 8?

OpenStudy (amistre64):

no, (-1)^2 = 1 and -8/1 = -8

OpenStudy (anonymous):

surely -8

OpenStudy (anonymous):

my mistake, -8 for sure lol. Thank you amistre64!

OpenStudy (amistre64):

;) good luck, its the algebra thats a killer lol